OpenStudy (anonymous):

a boats crew rowed 20 miles downstream , with the current , in 2 hours. the return trip upstream , against the current , covered the same distance , but took 5 hours. the crew rowing rate in still water is ? miles per hour the rate of the current is ? miles per hour

OpenStudy (anonymous):

Let the crew's rowing rate in still water be x miles per hour and that of stream be y mps. Then, time taken by crew rowing upstream = 20/(x+y) hrs. = 2 hrs. and time taken to row downstream = 20/(x-y) hrs. = 5 hrs. Two variables two equations solve them and get the answers

OpenStudy (anonymous):

the answer is 2 and 5 right ?

OpenStudy (anonymous):

i can not find the answer

OpenStudy (anonymous):

and i m on exam so can somebody help me